# On point spectrum of Jacobi matrices generated by iterations of quadratic polynomials

### v1

@@ A.1. Recursion for resolvent matrix functions
Iterating (A.1) we get
$W_{\varkappa}(z)=$ $-\frac{1}{z}\pi_{\varkappa}+\sum_{k=1}^{n-1}w^{k}_{0}(z){\mathfrak{P}}^{k-1}_{\varkappa}(\bar{z})^{*}\pi_{\hat{}\underline{s}^{k}{\varkappa}}{\mathfrak{P}}^{k-1}_{\varkappa}(z)$
@@
$+$ $\frac{(T^{\circ n})^{\prime}(z)}{2^{n}}({\mathfrak{P}}^{n-1}_{\varkappa}(\bar{z}))^{*}W_{\hat{}\underline{s}^{n}{\varkappa}}(T^{\circ n}(z)){\mathfrak{P}}^{n-1}_{\varkappa}(z)$ (A.5)
Since $W_{\varkappa}(z)$ is regular at the origin we get the following relation
$\pi_{\varkappa}=$ $\frac{1}{-T(0)}\Phi_{\varkappa}^{*}\pi_{\hat{\underline{s}}{\varkappa}}\Phi_{\varkappa}+\Phi_{\varkappa}^{*}\sum_{k=2}^{n-1}w^{k-1}_{0}(T(0)){\mathfrak{P}}^{k-2}_{\hat{\underline{s}}{\varkappa}}(T(0))^{*}\pi_{\hat{}\underline{s}^{k}{\varkappa}}{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}{\varkappa}}(T(0))\Phi_{\varkappa}$
$+$ $\Phi_{\varkappa}^{*}{\mathfrak{P}}^{n-2}_{\hat{\underline{s}}{\varkappa}}(T(0))^{*}\int w^{n-1}_{x}(T(0))\Sigma_{\hat{}\underline{s}^{n}{\varkappa}}(dx){\mathfrak{P}}^{n-2}_{\hat{\underline{s}}{\varkappa}}(T(0))\Phi_{\varkappa}$
where
$\Phi_{\varkappa}=\left.(z{\mathfrak{p}}_{\varkappa}(z))\right|_{z=0}=\Psi_{\varkappa}\Upsilon_{\varkappa}^{*}.$ (A.6)
Note that $\Upsilon_{\varkappa}\Upsilon_{\varkappa}^{*}=\pi_{\varkappa}$ and all summand in the RHS are positive. ∎
A.2. Periodic case
In this section we derive an explicit formula for eigenfunction of a matrix Ruelle operator if ${\varkappa}$ is periodic, i.e., there exists $d$ such that $\hat{\underline{s}}^{d}{\varkappa}={\varkappa}$.
The computations are based essentially on the identity (A.4). The measure $\Sigma_{\varkappa}(dx)$ is the eigen-measure for the adjoint to the Ruelle operator
$({\mathcal{M}}_{\varkappa}^{*}\cdots{\mathcal{M}}^{*}_{\hat{\underline{s}}^{d-1}{\varkappa}}\Sigma_{{\varkappa}})(dx)=\Sigma_{\varkappa}(dx).$
Of course the same is true for all shifts $\Sigma_{\hat{\underline{s}}^{k}{\varkappa}}(dx)$, $k=1,...,d-1$.
Having in mind Lemma 2.4, we are looking for the eigen-matrix-function $h_{\varkappa}(x)$,
$({\mathcal{M}}_{\hat{\underline{s}}^{d-1}{\varkappa}}\cdots{\mathcal{M}}_{\varkappa}h_{\varkappa})(x)=h_{\varkappa}(x),$ (A.7)
in the form
$h_{\varkappa}(x)=\frac{A_{1}}{x-T(0)}+\sum_{k=2}^{\infty}w_{x}^{k-1}(T(0))A_{k}.$
Due to (2.9),
$({\mathcal{M}}_{\varkappa}h_{\varkappa})(x)=$ $\frac{\Phi_{\varkappa}A_{1}\Phi_{{\varkappa}}^{*}}{-T(0)(x-T(0))}+\sum^{\infty}_{k=2}w_{0}^{k-1}(T(0))\frac{\Phi_{\varkappa}A_{k}\Phi_{{\varkappa}}^{*}}{x-T(0)}$
$+$ $\sum_{k=2}^{\infty}w_{x}^{k-1}(T(0)){\mathfrak{p}}_{\varkappa}(T^{\circ(k-1)}(0))A_{k-1}{\mathfrak{p}}_{\varkappa}(T^{\circ(k-1)}(0))^{*}.$
Similarly, applying ${\mathcal{M}}_{\hat{\underline{s}}{\varkappa}},\dots,{\mathcal{M}}_{\hat{\underline{s}}^{d-1}{\varkappa}}$, we obtain that (A.7) yields the recursion
$A_{k}={\mathfrak{p}}_{\hat{\underline{s}}^{d-1}{\varkappa}}(T^{\circ(k-1)}(0))\cdots{\mathfrak{p}}_{\varkappa}(T^{\circ(k-d)}(0))A_{k-d}$
×$\times\left({\mathfrak{p}}_{\hat{\underline{s}}^{d-1}{\varkappa}}(T^{\circ(k-1)}(0))\cdots{\mathfrak{p}}_{\varkappa}(T^{\circ(k-d)}(0))\right)^{*},\quad\text{for}\ k\geq d+1.$ (A.8)
Now, from first $d$ equations we can observe that $A_{1},\dots,A_{d}$ are necessarily of the form
$A_{1}=y_{1}\Psi_{\hat{\underline{s}}^{d-1}{\varkappa}}\Psi_{\hat{\underline{s}}^{d-1}{\varkappa}}^{*},\quad A_{2}=y_{2}{\mathfrak{p}}_{\hat{\underline{s}}^{d-1}{\varkappa}}(T(0))\Psi_{\hat{\underline{s}}^{d-2}{\varkappa}}\Psi_{\hat{\underline{s}}^{d-2}{\varkappa}}^{*}{\mathfrak{p}}_{\hat{\underline{s}}^{d-1}{\varkappa}}(T(0))^{*},$
and so on… till
$A_{d}=y_{d}{\mathfrak{p}}_{\hat{\underline{s}}^{d-1}{\varkappa}}(T^{\circ(d-1)}(0))\cdots{\mathfrak{p}}_{\hat{\underline{s}}{\varkappa}}(T(0))\Psi_{{\varkappa}}\Psi_{{\varkappa}}^{*}$
×$\times\left({\mathfrak{p}}_{\hat{\underline{s}}^{d-1}{\varkappa}}(T^{\circ(d-1)}(0))\cdots{\mathfrak{p}}_{\hat{\underline{s}}{\varkappa}}(T(0))\right)^{*}$ (A.9)
with scalars $y_{1},\dots,y_{d}$. In combination with (A.2) we obtain, $l\geq 1$,
  • (1)
    for $k=ld+1$
    $A_{k}=y_{1}{\mathfrak{P}}^{k-2}_{{\varkappa}}(T(0))\Psi_{\hat{\underline{s}}^{d-1}{\varkappa}}\Psi_{\hat{\underline{s}}^{d-1}{\varkappa}}^{*}{\mathfrak{P}}^{k-2}_{{\varkappa}}(T(0))^{*},$
  • (2)
    for $k=ld+2$
    $A_{k}=y_{2}{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}^{d-1}{\varkappa}}(T(0))\Psi_{\hat{\underline{s}}^{d-2}{\varkappa}}\Psi_{\hat{\underline{s}}^{d-2}{\varkappa}}^{*}{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}^{d-1}{\varkappa}}(T(0))^{*},$
  • (d)
    and so on… till $k=ld+d$
    $A_{k}=y_{d}{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}{\varkappa}}(T(0))\Psi_{{\varkappa}}\Psi_{{\varkappa}}^{*}{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}{\varkappa}}(T(0))^{*}.$
In this way we reduce (A.7) to a finite linear system
$\begin{pmatrix}y_{1}\\ \vdots\\ y_{d}\end{pmatrix}=L\begin{pmatrix}y_{1}\\ \vdots\\ y_{d}\end{pmatrix}=\begin{pmatrix}L_{11}&\ldots&L_{1d}\\ \vdots&&\vdots\\ L_{d1}&\ldots&L_{dd}\end{pmatrix}\begin{pmatrix}y_{1}\\ \vdots\\ y_{d}\end{pmatrix}$
with the matrix entries: for $j\not=d$
$L_{jn}=\sum_{k=ld+n-j}w_{0}^{k-1}(T(0))\langle{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}^{d-n+1}{\varkappa}}(T(0))\Psi_{\hat{\underline{s}}^{d-n}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{d-j}{\varkappa}}\rangle^{2},\ n\not=j+1,$
$L_{j(j+1)}=\frac{\langle\Psi_{\hat{\underline{s}}^{d-j-1}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{d-j}{\varkappa}}\rangle^{2}}{-T(0)}+\sum_{k=ld+1}w_{0}^{k-1}(T(0))\langle{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}^{d-j}{\varkappa}}(T(0))\Psi_{\hat{\underline{s}}^{d-j-1}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{d-j}{\varkappa}}\rangle^{2},$
@@ A.2. Periodic case
$L_{dn}=\sum_{k=ld+n}w_{0}^{k-1}(T(0))\langle{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}^{d-n+1}{\varkappa}}(T(0))\Psi_{\hat{\underline{s}}^{d-n}{\varkappa}},\Upsilon_{\varkappa}\rangle^{2},\quad n\geq 2,$
$L_{d1}=\frac{\langle\Psi_{\hat{\underline{s}}^{d-1}{\varkappa}},\Upsilon_{\varkappa}\rangle^{2}}{-T(0)}+\sum_{k=dl+1}w_{0}^{k-1}(T(0))\langle{\mathfrak{P}}^{k-2}_{{\varkappa}}(T(0))\Psi_{\hat{\underline{s}}^{d-1}{\varkappa}},\Upsilon_{\varkappa}\rangle^{2}.$
For a fixed $n$ we have
$\sum_{j=1}^{n}L_{jn}=$
$\frac{\langle\Psi_{\hat{\underline{s}}^{d-n}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{d-n+1}{\varkappa}}\rangle^{2}}{-T(0)}+\sum_{k=ld+1}w_{0}^{k-1}(T(0))\langle{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}^{d-n+1}{\varkappa}}(T(0))\Psi_{\hat{\underline{s}}^{d-n}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{d-n+1}{\varkappa}}\rangle^{2}$
$+\sum_{\begin{subarray}{c}k=ld+n-j\\ n\not=j+1\end{subarray}}w_{0}^{k-1}(T(0))\langle{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}^{d-n+1}{\varkappa}}(T(0))\Psi_{\hat{\underline{s}}^{d-n}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{d-j}{\varkappa}}\rangle^{2}$
$=\frac{\langle\Psi_{\hat{\underline{s}}^{d-n}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{d-n+1}{\varkappa}}\rangle^{2}}{-T(0)}+\sum_{k=2}^{\infty}w_{0}^{k-1}(T(0))\langle{\mathfrak{P}}^{k-2}_{\hat{\underline{s}}^{d-n+1}{\varkappa}}(T(0))\Psi_{\hat{\underline{s}}^{d-n}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{d-n+k}{\varkappa}}\rangle^{2}$
By (A.4) we obtain
$\sum_{j=1}^{n}L_{jk}=1.$
In other words we get a row-eigenvector $e$ of the matrix $L$,
$eL=e,\quad e=\begin{pmatrix}1&\ldots&1\end{pmatrix}.$
By the classical Perron-Frobenius Theorem $1$ is the spectral radius of the matrix $L$ and there is a unique column-eigenvector $f$, $Lf=f$, with positive entries $\{y_{j}\}_{1}^{d}$ normalized by the condition
$ef=\sum_{j=1}^{d}y_{j}=1.$
Thus we obtain the eigen-matrix-function $h_{\varkappa}(x)$, see (A.7), (A.2) and (1)–(d).
Remark A.3.
As a result we get an invariant measure,
$\mu(dx)=\text{\rm tr}(h_{\varkappa}(x)\Sigma_{\varkappa}(dx)),\quad\int g(T(x))\mu(dx)=\int g(x)\mu(dx).$
Thus we have a “computational” proof of the main theorem in the given particular case. It is easy to see that such a measure $\mu$ has no mass-points. Indeed, assume that $\mu(\{x_{0}\})=\delta>0$. Thus
$\mu(\{x_{0}\})=\sum_{T(y)=x_{0}}\mu(\{y\})$
If $x_{0}$ is not a fixed or periodic point preimages of different generations do not intersect, and we get that the total mass of them is infinite. If $x_{0}$ is periodic, then among preimages we have non-periodic $y_{0}$ and we apply the same arguments to this point. Note that the vectors
${\mathfrak{p}}_{\hat{\underline{s}}{\varkappa}}(T(0))\Psi_{{\varkappa}}=\begin{pmatrix}1&a_{\hat{\underline{s}}{\varkappa}}\\ 0&a_{\hat{\underline{s}}{\varkappa}+1}\end{pmatrix}\begin{pmatrix}1&0\\ 0&1/{T(0)}\end{pmatrix}\begin{pmatrix}a_{{\varkappa}-1}\\ a_{{\varkappa}}\end{pmatrix},\quad\Psi_{\hat{\underline{s}}{\varkappa}}=\begin{pmatrix}a_{\hat{\underline{s}}{\varkappa}}\\ a_{\hat{\underline{s}}{\varkappa}+1}\end{pmatrix}$
are linearly independent, therefore $h_{\varkappa}(x)>0$. Since $\mu(dx)$ does not contain mass-point components and $h_{\varkappa}(x)>0$ we get that $\Sigma_{\varkappa}(dx)$ also has no mass-point components.
A.3. Computations for $h_{k}(x)$ in a general case
We introduce $A\star B=ABA^{*}$. Then, see (A.2),
${\mathcal{M}}_{2^{n-k}{\varkappa}}\cdots{\mathcal{M}}_{2^{n-1}{\varkappa}}{\mathcal{M}}_{2^{n}{\varkappa}}w_{x}^{0}(T(0))I=w_{x}^{0}(T(0))f^{n-k}_{0}\Phi_{2^{n-k}{\varkappa}}\star I$
$+w_{x}^{1}(T(0))f^{n-k}_{1}({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T(0))\Phi_{2^{n-k+1}{\varkappa}})\star I+\dots$
$+w_{x}^{j}(T(0))f^{n-k}_{j}({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T^{\circ j}(0))\cdots{\mathfrak{p}}_{2^{n-k+j-1}{\varkappa}}(T(0))\Phi_{2^{n-k+j}{\varkappa}})\star I$
$+\dots+w_{x}^{k}(T(0))f^{n-k}_{k}({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T^{\circ k}(0))\cdots{\mathfrak{p}}_{2^{n-1}{\varkappa}}(T(0))\Phi_{2^{n}{\varkappa}})\star I$
$+w_{x}^{k+1}(T(0))({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T^{\circ(k+1)}(0)){\mathfrak{p}}_{2^{n-1}{\varkappa}}(T^{\circ 2}(0)){\mathfrak{p}}_{2^{n}{\varkappa}}(T(0)))\star I$
with certain constants $f^{n-k}_{j}$, $0\leq j\leq k$. Since, due to (2.5),
${\mathcal{M}}_{2^{n-(k+1)}{\varkappa}}w_{x}^{j}(T(0))A=$ $w_{0}^{j}(T(0))w_{x}^{0}(T(0))\Phi_{2^{n-(k+1)}{\varkappa}}\star A$
$+$ $w_{x}^{j+1}(T(0)){\mathfrak{p}}_{2^{n-(k+1)}{\varkappa}}(T^{\circ(j+1)}(0))\star A,$
these constants obey the following recurrence relations:
$f_{j+1}^{n-(k+1)}=f_{j}^{n-k}$
and
$f_{0}^{n-(k+1)}=w_{0}^{k+1}(T(0))\|({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T^{\circ(k+1)}(0))\cdots{\mathfrak{p}}_{2^{n}{\varkappa}}(T(0)))^{*}\Upsilon\|^{2}$
$+\sum_{j=0}^{k}w_{0}^{j}(T(0))\|({\mathfrak{p}}_{2^{n-k}{\varkappa}}(T^{\circ j}(0))\cdots{\mathfrak{p}}_{2^{n-k+j-1}{\varkappa}}(T(0))\Phi_{2^{n-k+j}{\varkappa}})^{*}\Upsilon\|^{2}f_{j}^{n-k}.$

### v2

@@ 4.2. Equivalence Lemma
We will use the following relations for $a_{\varkappa}^{2}$
$a_{2{\varkappa}}^{2}+a_{2{\varkappa}+1}^{2}$ $=\lambda,$
$a_{2{\varkappa}+1}^{2}a_{2{\varkappa}+2}^{2}$ $=a_{{\varkappa}+1}^{2},$
see [6, 24]. It yields an evident upper bound for $a_{\varkappa}$: $a_{\varkappa}^{2}\leq\lambda$ for all ${\varkappa}\in{\mathbb{Z}}_{2}$. It is known that $a_{{\varkappa}}\leq 1$ for an arbitrary even ${\varkappa}\in{\mathbb{Z}}_{2}$ [6]. It yields a lower bound for odd $\varkappa$}: $a^{2}_{{\varkappa}}\geq\lambda-1>1$. We will show that for ${\varkappa}\in{\mathcal{F}}_{N}$ coefficients $a_{{\varkappa}}$ are uniformly bounded from below
$a_{{\varkappa}}^{2}\geq C_{3}=\frac{\lambda-1}{\lambda^{N}},\quad\lambda>2.$ (4.4)
Using the above estimates, the proof follows by induction. Assume that $\varkappa$$^{\prime}$ is odd. As the base of induction we have
$a^{2}_{2{\varkappa}^{\prime}}=\frac{a^{2}_{{\varkappa}^{\prime}}}{a^{2}_{2{\varkappa}^{\prime}-1}}\geq\frac{\lambda-1}{\lambda}.$
Then we use induction
$a^{2}_{2^{n+1}{\varkappa}^{\prime}}=\frac{a^{2}_{2^{n}{\varkappa}^{\prime}}}{a^{2}_{2^{n+1}{\varkappa}^{\prime}-1}}\geq\frac{\lambda-1}{\lambda^{n}}\frac{1}{\lambda}=\frac{\lambda-1}{\lambda^{n+1}}.$ (4.5)
Since an arbitrary element of ${\mathcal{F}}_{N}$ contains at most $N$ consecutive zeros in its symbolic representation we have that ${\varkappa}=2^{n}{\varkappa}^{\prime}$, were $0\leq n\leq N$ and $\varkappa$$^{\prime}$ is odd. Thus (4.5) implies (4.4).
(ii) $\Rightarrow$ (i). Since for $x\in[-\xi,\xi]$, all terms in the sum (4.1) are positive we can estimate $h_{n}(x)$ by two last summands
$h_{n}(x)$ $\geq f^{n}_{0}({\varkappa})w^{0}_{x}(T(0))\Psi_{\hat{}\underline{s}^{n-1}{\varkappa}}\Psi_{\hat{}\underline{s}^{n-1}{\varkappa}}^{*}$
$\;\;+f^{n-1}_{0}({\varkappa})w^{1}_{x}(T(0)){\mathfrak{p}}_{\hat{}\underline{s}^{n-1}{\varkappa}}(T(0))\Psi_{\hat{}\underline{s}^{n-2}{\varkappa}}({\mathfrak{p}}_{\hat{}\underline{s}^{n-1}{\varkappa}}(T(0))\Psi_{\hat{}\underline{s}^{n-2}{\varkappa}})^{*}$
$\geq C_{2}\inf_{x\in[-\xi,\xi]}\{w^{0}_{x}(T(0)),w^{1}_{x}(T(0))\}$
$\;\;\times\left(\Psi_{\hat{}\underline{s}^{n-1}{\varkappa}}\Psi_{\hat{}\underline{s}^{n-1}{\varkappa}}^{*}+{\mathfrak{p}}_{\hat{}\underline{s}^{n-1}{\varkappa}}(T(0))\Psi_{\hat{}\underline{s}^{n-2}{\varkappa}}({\mathfrak{p}}_{\hat{}\underline{s}^{n-1}{\varkappa}}(T(0))\Psi_{\hat{}\underline{s}^{n-2}{\varkappa}})^{*}\right)$
Using the lower bound in (3.8), this implies
@@
$h_{n}(x)\geq\frac{C_{2}}{\lambda+\xi}\left(\Psi_{\hat{}\underline{s}{\varkappa}^{\prime}}\Psi_{\hat{}\underline{s}{\varkappa}^{\prime}}^{*}+{\mathfrak{p}}_{\hat{}\underline{s}{\varkappa}^{\prime}}(T(0))\Psi_{{\varkappa}^{\prime}}({\mathfrak{p}}_{\hat{}\underline{s}{\varkappa}^{\prime}}(T(0))\Psi_{{\varkappa}^{\prime}})^{*}\right),$ (4.6)
where ${\varkappa}^{\prime}=\hat{\underline{s}}^{n-2}{\varkappa}$.
Now we note that for two dimensional vectors $a$ and $b$
$\det(aa^{*}+bb^{*})={\det}^{2}\begin{pmatrix}a&b\end{pmatrix}$,
and therefore
$aa^{*}+bb^{*}\geq\frac{\det^{2}\begin{pmatrix}a&b\end{pmatrix}}{\|a\|^{2}+\|b\|^{2}}I.$
For the vectors
$a=\Psi_{\hat{}\underline{s}{\varkappa}^{\prime}}\quad\text{and}\quad b={\mathfrak{p}}_{\hat{}\underline{s}{\varkappa}^{\prime}}(T(0))\Psi_{{\varkappa}^{\prime}},$
by (4.4), we have
${\det}^{2}\begin{pmatrix}a&b\end{pmatrix}=a^{2}_{\hat{}\underline{s}{\varkappa}^{\prime}+{\epsilon}_{1}}a^{2}_{{\varkappa}^{\prime}+{\epsilon}_{2}}\geq C_{3}^{2},\quad{\epsilon}_{1},{\epsilon}_{2}\in\{-1,0,1\}.$
For the norms of these vectors we have
$\|a\|^{2}=\|\Psi_{\hat{\underline{s}}{\varkappa}^{\prime}}\|^{2}\leq 2\lambda$
@@
$\|b\|^{2}\leq\left(1+\frac{\|\Psi_{\hat{\underline{s}}{\varkappa}^{\prime}}\|^{2}}{T(0)^{2}}\right)\|\Psi_{{\varkappa}^{\prime}}\|^{2}\leq\left(1+\frac{2}{\lambda}\right)2\lambda.$
Combining all these estimates with (4.4) we get (i) with
$C_{1}=\frac{C_{2}C_{3}^{2}}{4(\lambda+\xi)(\lambda+1)}.$
(i) $\Rightarrow$ (ii). By (4.1), we have
$\langle h_{n}(0)\Upsilon_{\hat{\underline{s}}^{n}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{n}{\varkappa}}\rangle$ $=w^{n}_{0}(T(0))\langle{\mathfrak{P}}^{n}_{{\varkappa}}(T(0))^{*}\Upsilon_{\hat{\underline{s}}^{n}{\varkappa}},{\mathfrak{P}}^{n}_{{\varkappa}}(T(0))^{*}\Upsilon_{\hat{\underline{s}}^{n}{\varkappa}}\rangle$
$\;\;+\sum_{i=1}^{n}f^{i}_{0}({\varkappa})w^{n-i}_{0}(T(0))\langle{\mathfrak{P}}^{n-i}_{\hat{}\underline{s}^{i}{\varkappa}}(T(0))\Psi_{\hat{}\underline{s}^{i-1}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{n}{\varkappa}}\rangle^{2}.$ (4.7)
Due to (4.3)
$f^{n+1}_{0}({\varkappa})=w_{0}^{n}(T(0))\langle({\mathfrak{P}}^{n}_{{\varkappa}}(T(0)))^{\star 2}\Upsilon_{\hat{\underline{s}}^{n}{\varkappa}},\Upsilon_{\hat{\underline{s}}^{n}{\varkappa}}\rangle$
$+\sum_{i=1}^{n}{\mathfrak{L}}_{n+1,i}({\varkappa})f^{i}_{0}({\varkappa}).$ (4.8)